OpenStudy (anonymous):

HELP! Calculate the p value of each ion indicated: Ba2+, Mn2+, Cl- in a solution of BaCl2 7.65x10-3 M and MnCl2 1.54 M

OpenStudy (anonymous):

whats p vale..?

OpenStudy (anonymous):

I think is like the pH, but for every ion, pBa, pK, pCl...

OpenStudy (anonymous):

listen u just find normality after finding equivalent wieght of Ba+2,Mn+2,Cl- then put in the formula concnt of ion/concn of salt ok do ths much

OpenStudy (anonymous):

nw just put values pH=pKa + log concn of ion/concn of salt and then u ll get ans :) but yes in Ba case u ll get pOH dont forgot to subtract it by 14 as pH+ pOH=14

OpenStudy (anonymous):

equivalent weight /molecular weight=normality ok

OpenStudy (anonymous):

vicky u cal. normality

OpenStudy (anonymous):

27/208=normality 12 x10^-2 approx=normality nw Ka=14.3 x 10^-3 pKa=log 14.3 x10^-3 pKa=3-log14

OpenStudy (anonymous):

now after gettin pKa put it in the formula pH=pKa + log concn of ion/concn of salt u ll get it ok